OpenStudy (anonymous):

The speed of a sail-boat or other craft that does not plane is limited by the wave it makes – it can’t climb uphill over the front of the wave. What is the maximum speed you’d expect? Hint: relevant quantities might be the length l of the boat, the density ρ of the water, and the gravitational acceleration g. vboat = (L^X)(p^Y)(g^Z)

OpenStudy (anonymous):

do the dimensional analysis: \[[LT ^{-1}] = L ^{x-3y+z}M ^{y}T ^{-2z}\] comparing the powers will give : x=1/2 y=0 z=1/2